# ParallelTime: Dynamically Weighting the Balance of Short- and Long-Term Temporal Dependencies

### v1

@@ 6.2 Model Efficiency Analysis
Table 2 presents a comparison of MSE and MAE, Floating-Point Operations (FLOPs) for both training and inference and number of parameters, of our model against PatchTST across various prediction lengths using the Traffic dataset. The results show that our model requires significantly fewer FLOPs for both training and inference, achieves higher accuracy, and scales better with larger prediction lengths. This efficiency makes our model particularly well-suited for real-time long-term forecasting applications, where computational resources and speed are critical. For results on additional datasets, refer to Appendix 7.
@@ 7 Conclusion and Future Work
A key limitation of this study is that, due to limited GPU resources, we could not scale our model to include a large number of layers, which constrained our ability to train on larger datasets or more diverse tasks. Future research can explore the model’s potential as a foundation for time series analysis with minimal adjustments. Specifically, efforts can focus on fine-tuning the model for diverse tasks, such as anomaly detection, classification, and multi-step forecasting, across various domains.

### v2

@@ 6.2 Model Efficiency Analysis
Table 2 presents a comparison of MSE and MAE, Floating-Point Operations (FLOPs), and number of parameters, of our model against PatchTST across various prediction lengths using the Traffic dataset. The results show that our model requires significantly fewer FLOPs for both training and inference, achieves higher accuracy, and scales better with larger prediction lengths. This efficiency makes our model particularly well-suited for real-time long-term forecasting applications, where computational resources and speed are critical. For results on additional datasets, refer to Appendix 7.
@@ 7 Conclusion and Future Work
Future research can explore the model’s potential as a foundation for time series analysis with minimal adjustments. Specifically, efforts can focus on fine-tuning the model for diverse tasks, such as anomaly detection, classification, and multi-step forecasting, across various domains.